OpenStudy (anonymous):

A certain fraction is equivalent to 2/3. If 3 is added to the numerator, and 1 is added to the denominator, the resulting fraction is equivalent to 3/4. What is the numerator of the original fraction?

OpenStudy (mathstudent55):

x/y = 2/3 (x + 3)/(y + 1) = 3/4

OpenStudy (mathstudent55):

x is what you're looking for.

OpenStudy (anonymous):

i'm getting x=0 O.O

OpenStudy (anonymous):

No....

OpenStudy (anonymous):

mathstudent is correct with his equation. can you solve the system?

OpenStudy (anonymous):

I'll try again :P

OpenStudy (anonymous):

kk. :) .

OpenStudy (mathstudent55):

Since all you need is x, solve one equation for y and substitute into y of the other equation. This way you get an equation in x alone.